# Kernel Spectral Joint Embeddings for High-Dimensional Noisy Datasets using Duo-Landmark Integral Operators

### v2

@@
Xiucai Ding and Rong Ma Xiucai Ding is Associate Professor at Department of Statistics, University of California, Davis, CA 95616 (E-mail: xcading@ucdavis.edu). He is currently partially supported by NSF DMS-2515104.Rong Ma is Assistant Professor at Department of Biostatistics, T.H. Chan School of Public Health, Harvard University, Department of Data Science, Dana-Farber Cancer Institute, and the Broad Institute of MIT and Harvard, Boston, MA 02115 (E-mail: rongma@hsph.harvard.edu).
@@ 1.2 Our algorithm, main results and contributions
Our proposed method is summarized in Algorithm 1. The algorithm begins by assessing the information sharing or “alignability” between the two datasets of interest, avoiding unwanted distortions and misinterpretations that can arise from integrating unrelated data. The core of the algorithm involves linking data points from the point clouds $\mathcal{X}=\{\mathbf{x}_{i}\}_{1\leq i\leq n_{1}}$, $\mathcal{Y}=\{\mathbf{y}_{j}\}_{1\leq j\leq n_{2}}\subset\mathbb{R}^{p}$ via an asymmetric rectangular kernel affinity matrix, as described in (2). Importantly, we exclude “self-connections” within the individual dataset, $\mathcal{X}$ and $\mathcal{Y}$, and instead focus exclusively on establishing connections between the two datasets. This design choice is motivated by both theoretical and practical considerations, as detailed below.
@@
We finish this section by introducing some notation. To streamline our statements, we use the notion of stochastic domination, which is commonly adopted in random matrix theory (Erdős and Yau 2017) to syntactically simplify precise statements of the form “$\mathsf{X}^{(n)}$ is bounded with high probability by $\mathsf{Y}^{(n)}$ up to small powers of $n$.” More notation is given in the supplement.
@@ 1. Alignability screening:
$l(\widehat{\mathbf{z}}_{i})=\begin{cases}0,&\text{if }\mathbf{z}_{i}\in\{\mathbf{x}_{i}\},\\ 1,&\text{if }\mathbf{z}_{i}\in\{\mathbf{y}_{i}\}.\end{cases}$
@@ 3.1 Joint manifolds model, common structures and duo-landmark integral operators
For $\mathbf{x}_{i}^{0},1\leq i\leq n_{1},$ we assume that they are centered and i.i.d. sampled from some sub-Gaussian random vector $X:\Omega_{1}\rightarrow\mathbb{R}^{p}$ with respect to some probability space $(\Omega_{1},\mathcal{F}_{1},\mathbb{P}_{1}).$ Moreover, we assume that the range of $X$ is supported on an $m_{1}$-dimensional connected Riemannian manifold $\mathcal{M}_{1}$ isometrically embedded in $\mathbb{R}^{p}$ via $\iota_{1}:\mathcal{M}_{1}\rightarrow\mathbb{R}^{p}.$ We suppose that $\operatorname{dim}(\iota_{1}(\mathcal{M}_{1}))=r_{1}\leq p.$ Let ~$\widetilde{\mathcal{F}}_{1}$ be the Borel sigma algebra of $\iota_{1}(\mathcal{M}_{1})\subseteq\mathbb{R}^{p}$ and denotes ~$\widetilde{\mathbb{P}}_{1}$ be the probability measure of $X$ defined on ~$\widetilde{\mathcal{F}}_{1}$ induced from $\mathbb{P}_{1}.$ We assume that ~$\widetilde{\mathbb{P}}_{1}$ is absolutely continuous with respect to the volume measure on $\iota_{1}(\mathcal{M}_{1})$ whose density function is strictly positive. In addition, we assume that similar conditions hold for $\mathbf{y}_{j},1\leq j\leq n_{2},$ by replacing the quantities $(\Omega_{1},\mathcal{F}_{1},\mathbb{P}_{1},X,\mathcal{M}_{1},\iota_{1},m_{1},r_{1})$ by some other combinations $(\Omega_{2},\mathcal{F}_{2},\mathbb{P}_{2},Y,\mathcal{M}_{2},\iota_{2},m_{2},r_{2}).$ Finally, we assume that $\{\mathbf{x}_{i}^{0}\}$ and $\{\mathbf{y}_{j}^{0}\}$ are independent.
@@
Due to independence, we have that for $Z=X-Y,$ $\operatorname{Cov}(Z):=\mathbb{\Sigma}=\mathbb{\Sigma}_{1}+\mathbb{\Sigma}_{2}.$ Under Assumption 3.1, it follows that the rank $r=\operatorname{rank}(\mathbb{\Sigma})$ satisfies $\max\{r_{1},r_{2}\}\leq r\leq r_{1}+r_{2}$, and that there exists a rotation matrix $\mathbb{R}\in\mathbb{R}^{p\times p}$ such that $\mathbb{R}Z=(Z_{1},Z_{2},\cdots,Z_{r},0,\cdots,0)^{\top}\in\mathbb{R}^{p}.$
@@
Intuitively, the parameters $\{\theta_{1},...,\theta_{r}\}$ are jointly determined by the probability measures ~$\widetilde{\mathbb{P}}_{1}$ and ~$\widetilde{\mathbb{P}}_{2}$ capturing the signal structures of the individual datasets, as well as the structural sharing between the two embedded manifolds $\iota_{1}(\mathcal{M}_{1})$ and $\iota_{2}(\mathcal{M}_{2})$. Denote the sets
@@
We define the probability measure on $\mathcal{S}_{1}$ for $\mathbb{O}\mathbb{R}X$ based on ~$\widetilde{\mathbb{P}}_{1}$ as $\widetilde{\mathsf{P}}_{1},$ and define the measure on $\mathcal{S}_{2}$ for $\mathbb{O}\mathbb{R}Y$ based on ~$\widetilde{\mathbb{P}}_{2}$ as $\widetilde{\mathsf{P}}_{2}.$ For random vectors $X^{\prime}$ and $Y^{\prime}$ respectively defined on $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ associated with $\widetilde{\mathsf{P}}_{1}$ and $\widetilde{\mathsf{P}}_{2}$, according to Assumption 3.2, it is easy to see that there exists some $d^{\prime}$-dimensional ($d^{\prime}\leq r$) subvectors $X_{d}^{\prime}$ and $Y_{d}^{\prime},$ and some differential mapping $\vartheta^{\prime}:\mathbb{R}^{d^{\prime}}\rightarrow\mathbb{R}^{d^{\prime}}$ satisfying
@@
Suppose the assumptions of Proposition 3.4 hold, for $\ell\in\{1,2\}$ and $g_{\ell}\in\mathcal{L}_{2}(\mathcal{S}_{\ell},\widetilde{\mathsf{P}}_{\ell}),$ we define the duo-landmark integral operators $\mathcal{K}_{\ell}$ over $\mathcal{L}_{2}(\mathcal{S}_{\ell},\widetilde{\mathsf{P}}_{\ell})$ as
@@
$\mathcal{G}_{\ell}g_{\ell}(\bm{x}):=\int_{\mathcal{S}_{\ell}}k(\bm{x},\bm{y})g_{\ell}(\bm{y})\widetilde{\mathsf{P}}_{\ell}(\mathrm{d}\bm{y}),\qquad g_{\ell}\in\mathcal{L}_{2}(\mathcal{S}_{\ell},\widetilde{\mathsf{P}}_{\ell}).$ (16)
@@ 3.3.1 Robustness analysis
$\widehat{\phi}_{i}(\bm{x})=\frac{1}{\mu_{i}\sqrt{n_{1}}}\sum_{j=1}^{n_{1}}\widehat{k}_{1}(\bm{x},\mathbf{x}_{j})u_{ij},\ \widehat{\psi}_{i}(\bm{y})=\frac{1}{\mu_{i}\sqrt{n_{2}}}\sum_{j=1}^{n_{2}}\widehat{k}_{2}(\bm{y},\mathbf{y}_{j})v_{ij},$
@@ 4.1 Simultaneous clustering
Setting 1. The data $\mathbf{x}_{i},1\leq i\leq n_{1},$ are generated from (1), where the signals $\mathbf{x}_{i}^{0}$ are drawn from (43) with $K=6,\ \pi_{j}\equiv 1/6,\ \bm{\mu}_{j}=15\mathbf{e}_{j},\ \Sigma_{j}\equiv 9{\bf I}_{p},1\leq j\leq 6,$ where $\{\mathbf{e}_{j}\}$ are the standard Euclidean basis in $\mathbb{R}^{p}$, and the noises $\bm{\xi}_{i}$ are generated from multivariate Gaussian satisfying (27) with $\sigma_{1}^{2}=0.25$. The data $\mathbf{y}_{i},1\leq i\leq n_{2},$ are generated from (1), where the noises $\bm{\zeta}_{i}$ are multivariate Gaussian vectors satisfying (27) with $\sigma_{2}^{2}=1$, and the signals $\mathbf{y}_{i}^{0}$ satisfy $\mathbf{y}_{i}^{0}=\mathbf{z}^{0}_{i}+\mathbb{w}^{0}_{i}$, with $\mathbf{z}^{0}_{i}$ generated from (43) with $K=6,\ \pi_{j}\equiv 1/6,\ \bm{\mu}_{j}=15\mathbf{e}_{j},\ \Sigma_{j}\equiv 9{\bf I}_{p},1\leq j\leq 6$, and $\mathbb{w}^{0}_{i}$ having all their components being zero except for its 6-th to 25-th components which are independently drawn from a uniform distribution between $[-3\tau,\tau]$. In this way, both datasets contain a low-dimensional Gaussian mixture cluster structure with the common cluster centers, whereas the second dataset $\{\mathbf{y}_{i}\}_{1\leq i\leq n_{2}}$ is much noisier and contains an additional data-specific signal structure in the subspace nearly orthogonal to the cluster centers $\{\bm{\mu}_{j}\}$.
@@ 4.2 Nonlinear manifold learning
In the second study, we consider enhancing the embeddings of low-dimensional manifold structures contained in a high-dimensional noisy dataset $\{\mathbf{y}_{j}\}_{1\leq j\leq n_{2}}$, with the help of an external dataset $\{\mathbf{x}_{i}\}_{1\leq i\leq n_{1}}$ which contains stronger and cleaner signals. For the data point $\mathbf{y}_{j}$ from the nosier dataset, in terms of the model setup (1) and noise setup (27), we let $\bm{\zeta}_{j}$ be generated from multivariate Gaussian distribution with $\sigma_{2}^{2}=1.$ For the clean signal part $\mathbf{y}_{j}^{0},$ its first three coordinates are uniformly drawn from a torus in $\mathbb{R}^{3}$, i.e., for $u,v\in[0,2\pi)$, $\mathbf{y}_{j1}^{0}=\theta(2+0.8\cos u)\cos v,\mathbf{y}_{j2}^{0}=\theta(2+0.8\cos u)\sin v$ and $\mathbf{y}_{j3}^{0}=0.8\theta\sin u,$ where $\theta=0.2n^{1/2}$ characterizing the overall signal strength of the torus. For the subsequent coordinates of $\mathbf{y}_{i}^{0},$ the first 20 components (i.e., the 4-th to 23-th components) are independently drawn from a uniform distribution on $[-8,8]$, whereas the rest components are zero. The dataset $\{\mathbf{y}^{0}_{j}\}_{1\leq j\leq n_{2}}$ contains two mutually orthogonal low-dimensional signal structures. For the data point $\mathbf{x}_{i}$ from the cleaner dataset, in terms of the model setup (1) and noise setup (27), we generate $\bm{\xi}_{i}$ from multivariate Gaussian distribution with $\sigma_{1}^{2}=0.16.$ For the clean signal part $\mathbf{x}_{j}^{0},$ its first three coordinates are uniformly drawn from the same torus in $\mathbb{R}^{3}$, and the rest of the coordinates are set as zeros. As such, $\{\mathbf{x}_{i}\}_{1\leq i\leq n_{1}}$ is less noisy and only contains partially shared signal structure with $\{\mathbf{y}_{j}\}_{1\leq i\leq n_{2}}$, i.e., the torus structure.
@@ 5 Applications in integrative single-cell omics analysis
Throughout the paper, the following conventions and notations will be consistently used. We interchangeably use the notation $\mathsf{X}=\mathrm{O}_{\prec}(\mathsf{Y})$, $\mathsf{X}\prec\mathsf{Y}$ or $\mathsf{Y}\succ\mathsf{X}$ if $\mathsf{X}$ is stochastically dominated by $\mathsf{Y}$, uniformly in $u\in\mathsf{U}^{(n)}$, when there is no risk of confusion. We say that an $n$-dependent event $\Psi\equiv\Psi(n)$ holds with high probability if for any large $D>1$, there exists $n_{0}=n_{0}(D)>0$ so that $\mathbb{P}(\Psi)\geq 1-n^{-D},$ for all $n\geq n_{0}.$ For two sequences of deterministic positive values $\{a_{n}\}$ and $\{b_{n}\},$ we write $a_{n}=\mathrm{O}(b_{n})$ if $a_{n}\leq Cb_{n}$ for some positive constant $C>0.$ In addition, if both $a_{n}=\mathrm{O}(b_{n})$ and $b_{n}=\mathrm{O}(a_{n}),$ we write $a_{n}\asymp b_{n}.$ Moreover, we write $a_{n}=\mathrm{o}(b_{n})$ if $a_{n}\leq c_{n}b_{n}$ for some positive sequence $c_{n}\to 0.$ For any probability measure $\mathsf{P}$ over $\Omega$, we denote $\mathcal{L}_{2}(\Omega,\mathsf{P})$ as the collection of $L_{2}$-integrable functions with respect to $\mathsf{P}$, that is, for any $f\in\mathcal{L}_{2}(\Omega,\mathsf{P})$, we have $\|f\|_{\mathsf{P}}=\sqrt{\int_{\Omega}|f(y)|^{2}\mathsf{P}(dy)}<\infty$. For a vector $\mathbb{a}=(a_{1},...,a_{n})^{\top}\in\mathbb{R}^{n}$, we define its $\ell_{p}$ norm as $\|\mathbb{a}\|_{p}=\big{(}\sum_{i=1}^{n}|a_{i}|^{p}\big{)}^{1/p}$. We denote $\text{diag}(a_{1},...,a_{n})\in\mathbb{R}^{n\times n}$ as the diagonal matrix whose $i$-th diagonal entry is $a_{i}$. For a matrix $\mathbb{A}=(a_{ij})\in\mathbb{R}^{n\times n}$, we define its Frobenius norm as $\|\mathbb{A}\|_{F}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}a^{2}_{ij}}$, and its operator norm as $\|\mathbb{A}\|=\sup_{\|\mathbb{x}\|_{2}\leq 1}\|\mathbb{A}\mathbb{x}\|_{2}$. For any integer $n>0$, we denote the set $[n]=\{1,2,...,n\}$. For a random vector $\mathbf{g},$ we say it is sub-Gaussian if $\mathbb{E}\exp(\mathbf{a}^{\top}\mathbf{g})\leq\exp\left(\|\mathbf{a}\|_{2}^{2}/2\right)$ for any deterministic vector $\mathbf{a}.$ Throughout, $C,C_{1},C_{2},...$ are universal constants independent of $n$, and can vary from line to line. We denote ${\bf I}_{n}$ as an $n$-dimensional identity matrix.

### v3

# arXiv:2405.12317v3[stat.ML] 27 Feb 2026

@@
Xiucai Ding and Rong Ma Xiucai Ding is Associate Professor at Department of Statistics, University of California, Davis, CA 95616 (E-mail: xcading@ucdavis.edu). He is currently partially supported by NSF DMS-2515104.Rong Ma is Assistant Professor at Department of Biostatistics, T.H. Chan School of Public Health, Harvard University, Department of Data Science, Dana-Farber Cancer Institute, and the Broad Institute of MIT and Harvard, Boston, MA 02115 (E-mail: rongma@hsph.harvard.edu). Rong gratefully acknowledges Boris Landa and Yuval Kluger for constructive discussions on methodology and numerical simulations during the development of the methods for a related problem in Landa et al. (2024).
@@ 1.2 Our algorithm, main results and contributions
Our proposed method is summarized in Algorithm 1. Influenced by Shen and Wu (2022), the algorithm begins by assessing the information sharing or “alignability” between the two datasets of interest, avoiding unwanted distortions and misinterpretations that can arise from integrating unrelated data. The core of the algorithm involves linking data points from the point clouds $\mathcal{X}=\{\mathbf{x}_{i}\}_{1\leq i\leq n_{1}}$, $\mathcal{Y}=\{\mathbf{y}_{j}\}_{1\leq j\leq n_{2}}\subset\mathbb{R}^{p}$ via an asymmetric rectangular kernel affinity matrix, as described in (2). Importantly, we exclude “self-connections” within the individual dataset, $\mathcal{X}$ and $\mathcal{Y}$, and instead focus exclusively on establishing connections between the two datasets. This design choice is motivated by both theoretical and practical considerations, as detailed below.
@@
This work was developed concurrently with Landa et al. (2024). At a high level, both papers study joint embeddings of two datasets using singular vectors/eigenfunctions from an asymmetric cross-dataset affinity matrix/operator. Several elements of the present work were indeed influenced by the ideas there. For example, the normalization of the asymmetric matrix and aspects of the numerical simulation design. While the two works address related problems, Landa et al. (2024) develop an optimal-transport-based approach grounded in a diffusion-type framework that accommodates unshared dataset-specific deformations and corruptions (e.g., shifts/dilations and nuisance structures).
@@
We finish this section by introducing some notation. To streamline our statements, we use the notion of stochastic domination, which is commonly adopted in random matrix theory (Erdős and Yau 2017) to syntactically simplify precise statements of the form $\mathsf{X}^{(n)}$ is bounded with high probability by $\mathsf{Y}^{(n)}$ up to small powers of nn.” More notation is given in the supplement.
@@ 1. Alignability screening:
[…]
@@ 3.1 Joint manifolds model, common structures and duo-landmark integral operators
For $\mathbf{x}_{i}^{0},1\leq i\leq n_{1},$ we assume that they are centered and i.i.d. sampled from some sub-Gaussian random vector $X:\Omega_{1}\rightarrow\mathbb{R}^{p}$ with respect to some probability space $(\Omega_{1},\mathcal{F}_{1},\mathbb{P}_{1}).$ Moreover, we assume that the range of $X$ is supported on an $m_{1}$-dimensional connected Riemannian manifold $\mathcal{M}_{1}$ isometrically embedded in $\mathbb{R}^{p}$ via $\iota_{1}:\mathcal{M}_{1}\rightarrow\mathbb{R}^{p}.$ We suppose that $\operatorname{dim}(\iota_{1}(\mathcal{M}_{1}))=r_{1}\leq p.$ Let $\widetilde{\mathcal{F}}_{1}$ be the Borel sigma algebra of $\iota_{1}(\mathcal{M}_{1})\subseteq\mathbb{R}^{p}$ and denotes $\widetilde{\mathbb{P}}_{1}$ be the probability measure of $X$ defined on $\widetilde{\mathcal{F}}_{1}$ induced from $\mathbb{P}_{1}.$ We assume that $\widetilde{\mathbb{P}}_{1}$ is absolutely continuous with respect to the volume measure on $\iota_{1}(\mathcal{M}_{1})$ whose density function is strictly positive. In addition, we assume that similar conditions hold for $\mathbf{y}_{j},1\leq j\leq n_{2},$ by replacing the quantities $(\Omega_{1},\mathcal{F}_{1},\mathbb{P}_{1},X,\mathcal{M}_{1},\iota_{1},m_{1},r_{1})$ by some other combinations $(\Omega_{2},\mathcal{F}_{2},\mathbb{P}_{2},Y,\mathcal{M}_{2},\iota_{2},m_{2},r_{2}).$ Finally, we assume that $\{\mathbf{x}_{i}^{0}\}$ and $\{\mathbf{y}_{j}^{0}\}$ are independent.
@@
Due to independence, we have that for Z=XY,Z=X-Y, $\operatorname{Cov}(Z):=\mathbb{\Sigma}=\mathbb{\Sigma}_{1}+\mathbb{\Sigma}_{2}.$ Under Assumption 3.1, it follows that the rank $r=\operatorname{rank}(\mathbb{\Sigma})$ satisfies $\max\{r_{1},r_{2}\}\leq r\leq r_{1}+r_{2}$, and that there exists a rotation matrix $\mathbb{R}\in\mathbb{R}^{p\times p}$ such that $\mathbb{R}Z=(Z_{1},Z_{2},\cdots,Z_{r},0,\cdots,0)^{\top}\in\mathbb{R}^{p}.$
@@
Intuitively, the parameters $\{\theta_{1},...,\theta_{r}\}$ are jointly determined by the probability measures $\widetilde{\mathbb{P}}_{1}$ and $\widetilde{\mathbb{P}}_{2}$ capturing the signal structures of the individual datasets, as well as the structural sharing between the two embedded manifolds $\iota_{1}(\mathcal{M}_{1})$ and $\iota_{2}(\mathcal{M}_{2})$. Denote the sets
@@
We define the probability measure on $\mathcal{S}_{1}$ for $\mathbb{O}\mathbb{R}X$ based on $\widetilde{\mathbb{P}}_{1}$ as $\widetilde{\mathsf{P}}_{1},$ and define the measure on $\mathcal{S}_{2}$ for $\mathbb{O}\mathbb{R}Y$ based on $\widetilde{\mathbb{P}}_{2}$ as $\widetilde{\mathsf{P}}_{2}.$ For random vectors $X^{\prime}$ and $Y^{\prime}$ respectively defined on $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ associated with $\widetilde{\mathsf{P}}_{1}$ and $\widetilde{\mathsf{P}}_{2}$, according to Assumption 3.2, it is easy to see that there exists some $d^{\prime}$-dimensional ($d^{\prime}\leq r$) subvectors $X_{d}^{\prime}$ and $Y_{d}^{\prime},$ and some differential mapping $\vartheta^{\prime}:\mathbb{R}^{d^{\prime}}\rightarrow\mathbb{R}^{d^{\prime}}$ satisfying
@@
[…]
@@
[…]
@@ 3.3.1 Robustness analysis
$\widehat{\phi}_{i}(\bm{x})=\frac{1}{\mu_{i}\sqrt{n_{1}}}\sum_{j=1}^{n_{1}}\widehat{k}_{1}(\bm{x},\mathbf{x}_{j})u_{ij},\ \widehat{\psi}_{i}(\bm{y})=\frac{1}{\mu_{i}\sqrt{n_{2}}}\sum_{j=1}^{n_{2}}\widehat{k}_{2}(\bm{y},\mathbf{y}_{j})v_{ij},\$
@@ 4.1 Simultaneous clustering
Setting 1. The data $\mathbf{x}_{i},1\leq i\leq n_{1},$ are generated from (1), where the signals $\mathbf{x}_{i}^{0}$ are drawn from (43) with $K=6,\ \pi_{j}\equiv 1/6,\ \bm{\mu}_{j}=15\mathbf{e}_{j},\ \Sigma_{j}\equiv 9{\bf I}_{p},1\leq j\leq 6,$ where $\{\mathbf{e}_{j}\}$ are the standard Euclidean basis in $\mathbb{R}^{p}$, and the noises $\bm{\xi}_{i}$ are generated from multivariate Gaussian satisfying (27) with $\sigma_{1}^{2}=0.25$. The data $\mathbf{y}_{i},1\leq i\leq n_{2},$ are generated from (1), where the noises $\bm{\zeta}_{i}$ are multivariate Gaussian vectors satisfying (27) with $\sigma_{2}^{2}=1$, and the signals $\mathbf{y}_{i}^{0}$ satisfy $\mathbf{y}_{i}^{0}=\mathbf{z}^{0}_{i}+\mathbb{w}^{0}_{i}$, with $\mathbf{z}^{0}_{i}$ generated from (43) with $K=6,\ \pi_{j}\equiv 1/6,\ \bm{\mu}_{j}=15\mathbf{e}_{j},\ \Sigma_{j}\equiv 9{\bf I}_{p},1\leq j\leq 6$, and $\mathbb{w}^{0}_{i}$ having all their components being zero except for its 6-th to 25-th components which are independently drawn from a uniform distribution between $3\tau,\tau]$. In this way, both datasets contain a low-dimensional Gaussian mixture cluster structure with the common cluster centers, whereas the second dataset $\{\mathbf{y}_{i}\}_{1\leq i\leq n_{2}}$ is much noisier and contains an additional data-specific signal structure in the subspace nearly orthogonal to the cluster centers $\{\bm{\mu}_{j}\}$.
@@ 4.2 Nonlinear manifold learning
In the second study, we consider enhancing the embeddings of low-dimensional manifold structures contained in a high-dimensional noisy dataset $\{\mathbf{y}_{j}\}_{1\leq j\leq n_{2}}$, with the help of an external dataset $\{\mathbf{x}_{i}\}_{1\leq i\leq n_{1}}$ which contains stronger and cleaner signals. For the data point $\mathbf{y}_{j}$ from the nosier dataset, in terms of the model setup (1) and noise setup (27), we let $\bm{\zeta}_{j}$ be generated from multivariate Gaussian distribution with $\sigma_{2}^{2}=1.$ For the clean signal part $\mathbf{y}_{j}^{0},$ its first three coordinates are uniformly drawn from a torus in $\mathbb{R}^{3}$, i.e., for $u,v\in[0,2\pi)$, $\mathbf{y}_{j1}^{0}=\theta(2+0.8\cos u)\cos v,\mathbf{y}_{j2}^{0}=\theta(2+0.8\cos u)\sin v$ and $\mathbf{y}_{j3}^{0}=0.8\theta\sin u,$ where $\theta=0.2n^{1/2}$ characterizing the overall signal strength of the torus. For the subsequent coordinates of $\mathbf{y}_{i}^{0},$ the first 20 components (i.e., the 4-th to 23-th components) are independently drawn from a uniform distribution on [8,8][-8,8], whereas the rest components are zero. The dataset $\{\mathbf{y}^{0}_{j}\}_{1\leq j\leq n_{2}}$ contains two mutually orthogonal low-dimensional signal structures. For the data point $\mathbf{x}_{i}$ from the cleaner dataset, in terms of the model setup (1) and noise setup (27), we generate $\bm{\xi}_{i}$ from multivariate Gaussian distribution with $\sigma_{1}^{2}=0.16.$ For the clean signal part $\mathbf{x}_{j}^{0},$ its first three coordinates are uniformly drawn from the same torus in $\mathbb{R}^{3}$, and the rest of the coordinates are set as zeros. As such, $\{\mathbf{x}_{i}\}_{1\leq i\leq n_{1}}$ is less noisy and only contains partially shared signal structure with $\{\mathbf{y}_{j}\}_{1\leq i\leq n_{2}}$, i.e., the torus structure.
@@ 5 Applications in integrative single-cell omics analysis
Throughout the paper, the following conventions and notations will be consistently used. We interchangeably use the notation $\mathsf{X}=\mathrm{O}_{\prec}(\mathsf{Y})$, $\mathsf{X}\prec\mathsf{Y}$ or $\mathsf{Y}\succ\mathsf{X}$ if $\mathsf{X}$ is stochastically dominated by $\mathsf{Y}$, uniformly in $u\in\mathsf{U}^{(n)}$, when there is no risk of confusion. We say that an $n$-dependent event $\Psi\equiv\Psi(n)$ holds with high probability if for any large $D>1$, there exists $n_{0}=n_{0}(D)>0$ so that $\mathbb{P}(\Psi)\geq 1-n^{-D},$ for all $n\geq n_{0}.$ For two sequences of deterministic positive values $\{a_{n}\}$ and $\{b_{n}\},$ we write $a_{n}=\mathrm{O}(b_{n})$ if $a_{n}\leq Cb_{n}$ for some positive constant $C>0.$ In addition, if both $a_{n}=\mathrm{O}(b_{n})$ and $b_{n}=\mathrm{O}(a_{n}),$ we write $a_{n}\asymp b_{n}.$ Moreover, we write $a_{n}=\mathrm{o}(b_{n})$ if $a_{n}\leq c_{n}b_{n}$ for some positive sequence $c_{n}\to 0.$ For any probability measure $\mathsf{P}$ over $\Omega$, we denote $\mathcal{L}_{2}(\Omega,\mathsf{P})$ as the collection of $L_{2}$-integrable functions with respect to $\mathsf{P}$, that is, for any $f\in\mathcal{L}_{2}(\Omega,\mathsf{P})$, we have $\|f\|_{\mathsf{P}}=\sqrt{\int_{\Omega}|f(y)|^{2}\mathsf{P}(dy)}<\infty$. For a vector $\mathbb{a}=(a_{1},...,a_{n})^{\top}\in\mathbb{R}^{n}$, we define its $\ell_{p}$ norm as $\|\mathbb{a}\|_{p}=\big(\sum_{i=1}^{n}|a_{i}|^{p}\big)^{1/p}$. We denote $\text{diag}(a_{1},...,a_{n})\in\mathbb{R}^{n\times n}$ as the diagonal matrix whose $i$-th diagonal entry is $a_{i}$. For a matrix $\mathbb{A}=(a_{ij})\in\mathbb{R}^{n\times n}$, we define its Frobenius norm as $\|\mathbb{A}\|_{F}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}a^{2}_{ij}}$, and its operator norm as $\|\mathbb{A}\|=\sup_{\|\mathbb{x}\|_{2}\leq 1}\|\mathbb{A}\mathbb{x}\|_{2}$. For any integer $n>0$, we denote the set $n]=\{1,2,...,n\}$. For a random vector $\mathbf{g},$ we say it is sub-Gaussian if $\mathbb{E}\exp(\mathbf{a}^{\top}\mathbf{g})\leq\exp\left(\|\mathbf{a}\|_{2}^{2}/2\right)$ for any deterministic vector $\mathbf{a}.$ Throughout, $C,C_{1},C_{2},$... are universal constants independent of $n$, and can vary from line to line. We denote ${\bf I}_{n}$ as an $n$-dimensional identity matrix.